What is the final dilution of a sample after 3 successive 1:10 dilutions and two 1:100 dilutions
Answer #1

decreases by the same factor in each successive step.

In serial dilutions, you multiply the dilution factors for each step.

The dilution factor or the dilution is the initial volume divided by the final volume.

DF=Vi/Fi

1/10 * 1/100 * 1/100 = 1/100000